Important industrial processes, such as converting iron ore to iron and then to steel, involve coupling a non-spontaneous reaction, such as reducing Fe2O3 (iron ore)to metallic iron, with a spontaneous process such as the oxidation of carbon in CO to CO2:

Fe2O3 (s) + 3 CO (g)  2 Fe (s) + 3 CO2 (g)

Reagent

ΔHºf (kJ/mol)

ΔSº(J/mol*K)

Fe2O3 (s)

-824.2

87.4

CO (g)

-110.5

197.7

Fe(s)

0

27.3

CO2 (g)

-393.5

213.8

  1. Use the thermodynamic data provided above to determine if this reaction is endothermic or exothermic under standard conditions (298 K). Show your work!
  2. Is the reaction spontaneous under standard conditions?  Show your work!
  3. Calculate the value of the equilibrium constant, Kp, at 298 K using the value of ΔGºrxn you obtained in part b. (If you did not find the value of ΔGºrxn in part b, you may assume ΔGºrxn = -25.0 kJ/mol.)
  4. If the reaction was performed under non-standard conditions at 200 ºC, using 6.0 g Fe2O3, 0.245 atm CO, and 0.175 atm CO2what is ΔGrxn and is the reaction spontaneous?

Show work please stuck on this homework problem
Answer #1

FG 03 (s) + 32010) -+ 2Fe (s) + 300219) Hnen = (4H) Product - (H) memutant = 2414)re(s) + 3 (04)60,ce) – (H)F6,0360) – 3 (Mencit25.0X103J/mol =48.314 Jimori ! x298 K enkp enke 25*103 8.314x298 enkp=10.09 [Rp = 2:41x104 / Any O G = = RT enkp+ RT en Q